RC Meeting and Run Plan, Tuesday December 19th 2019

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are at 4:15 PM, 2nd floor Counting House Meeting Room

  • Bluejeans remote connection information:
 To join the Meeting:
 To join via Room System:
 Video Conferencing System: bjn.vc -or-
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links

Yesterday's Run Plan

Wednesday Swing + Thursday Owl shifts:

  • Elastics data taken on reference cell filled with N2, H2, empty and finally He3. Online analysis plots available for the elastics data taken yesterday.
  • N2, H2, He3 pressure curves runs data taken
  • BCM calibration (+ parasitic MOLLER run) data taken
  • Could not perform the energy measurement due to arc2 energy lock limitations.
  • Beam charge asymmetry measurements performed. "Flipper" script ran after HWP retraction.
  • Data taken on Brianna!!
    • longitudinal asymmetries (1 hr DAQ running time)
    • Delta resonance asymmetry run (2 hrs with HWP + 2 hrs w/ HWP)

Target related work:

  • Cooling jets flow rate increased as they indicated temporary instability
  • Polarimetry debugging
    • NMR -- Target field rotated by 180 to checkout polarization. NMR performed at regular intervals. Shows some polarization that needs to be understood.
    • EPR -- working but noisy
    • pNMR -- turned off to be debugged later


  • Paddles that need to be turned off optimized to look at elastic peak.

Current Run Plan

Thursday -->Day shift (+ Swing + Owl):

  • Spectrometer rotation from 11.7 to 8.5 SHMS - done
  • Energy measurement - done
  • MOLLER setup - in progress
  • Optics delta scan runs with carbon foils (Jones et al will do it in the evening)
  • Back to He3 longitudinal asymmetry measurements till 8:00 AM.
  • Perform NMR at regular intervals
  • Day by Day Options [1]

Notices in Effect for Shift Crew

  • Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
  • Magnet policies
    • HMS Q1 must be ramped down in 50A steps in all cases
    • Do not ramp HMS Q2 to zero, only go to 10A and then switch off power.
  • Inform the RC of any access that the Target Team requires. The RC will coordinate with RadCon to make everything run smoothly.
  • The operators have been somewhat careful with our beam orbit locks. Please make sure that the orbit locks are enabled and the beam does not drift from its nominal position.
  • Ramp rate lock instructions are here and added to checklist[2]
  • Temporary Laser instructions can be found here [3]

Current Issues

  • Polarimetry debugging in progress

Ongoing Work in the Hall

Planned Accelerator Configuration Changes

Special Requests

  • Keep target expert schedule filled out on whiteboard